# LIFBench: Evaluating the Instruction Following Performance and Stability of Large Language Models in Long-Context Scenarios

### v2

@@ Abstract
Xiaodong Wu  Minhao Wang  Yichen Liu  Xiaoming Shi  He Yan §  Xiangju Lu §  Junmin Zhu §  Wei Zhang thanks: Corresponding author. East China Normal University  §iQIYI Inc {51255901079,51275901104,51275901148}@stu.ecnu.edu.cn  xmshi@ir.hit.edu.cn §{yanhe, luxiangju, zhujunmin}@qiyi.com  *zhangwei.thu2011@gmail.com
@@ 1 Introduction
As Large Language Models (LLMs) continue to make significant strides across practical applications Achiam et al. (2023); Chowdhery et al. (2023); Brown (2020), their performance in natural language processing (NLP) tasks has reached unprecedented levels. These tasks span text generation Que et al. (2024); Tan et al. (2024); Zhang et al. (2024b), complex reasoning Parmar et al. (2024), and problem-solving Lu et al. (2024); Li et al. (2024b). Despite these achievements, significant challenges remain. On the one hand, LLMs often struggle to accurately and consistently follow human instructions, such as restating input content precisely or adhering stably to specific formatting constraints He et al. (2024). On the other hand, studies show that as input length increases, the LLMs’ performance in tasks such as reasoning Levy et al. (2024), retrieval Li et al. (2024a), and general NLP Bai et al. (2023) deteriorates. These challenges pose substantial barriers to their effectiveness in real-world applications.
@@
For evaluation, traditional metrics for downstream tasks are often unsuitable for complex instruction-following scenarios Honovich et al. (2023). Moreover, while many studies rely on GPT-4 for automated and open-ended assessment, these approaches encounter limitations due to notable gaps between GPT-4 and human experts Qin et al. (2024b); Jiang et al. (2023), as well as potential bias problems Wang et al. (2023). To address these challenges, we propose LIFEval, a systematic and efficient evaluation method for complex LLM responses, without relying on LLMs or human evaluators. Specifically, by designing task-specific scoring rubrics, we decompose evaluations into fine-grained and quantifiable scoring points, each of which can be assessed automatically. In addition, through the score-capability map and a novel metric—IFS, LIFEval provides insights into models’ fundamental capabilities and stability from various perspectives.
@@ 2 Related Work
Several studies focus on LLMs’ performance in long contexts. These benchmarks collect data from traditional NLP tasks (e.g., summarization and question answering (QA)) to form comprehensive datasets containing long data Shaham et al. (2022, 2023); Dong et al. (2024); Li et al. (2024c). Additionally, due to the excellent performance of LLMs on open-ended tasks, some benchmarks have also designed synthetic tasks to better probe the models’ ability in math, reasoning, and logic Bai et al. (2023); Zhang et al. (2024c); Kwan et al. (2023); Wang et al. (2024); Hsieh et al. (2024); Li et al. (2024a). For evaluation, some works adopt regular metrics (e.g., Acc, ROUGE, and BLEU), which can be obtained through automated calculations. However, some open-ended tasks cannot be effectively evaluated using these metrics, hence powerful LLMs, such as GPT-4, are used as alternative evaluators. For example, the studies An et al. (2023); Li et al. (2023) feed the model predictions and ground-truth answers to the GPT-4 evaluator, which is tasked with conducting a direct scoring or comparison to evaluate baselines’ performance on partial tasks. In contrast to LIFBench, these benchmarks only assess problem-solving capabilities in long-context scenarios, overlooking the challenges of instruction following that arise in real-world applications.
Besides, some other efforts Jiang et al. (2023); Jing et al. (2023); Zhou et al. (2023a); Qin et al. (2024a) focus on diverse constraint types in real-world tasks, aiming to assess LLMs’ abilities in understanding and meeting these constraints. Given the complexities of evaluating instruction-following abilities, several studies introduce meaningful innovations in their evaluation methodologies to better tackle this multifaceted challenge. The studies Cook et al. (2024); Wen et al. (2024); Qin et al. (2024b); Zhang et al. (2024a) decompose instructions into checklists composed of YES/NO questions or PASS/FAIL criteria, which answers should meet. CELLO He et al. (2024) defines four criteria that should be assessed as they can encompass common errors made by LLMs, and develops automated evaluation metrics to reflect models’ ability in complex instruction-following scenarios. However, these studies fail to address model performance in long-context instruction-following scenarios and largely overlook stability in the instruction-following process.
  Scenario Task ID #Samples
List Single-ID LSI 180
Multi-ID LMI 150
Offset-ID LOI 396
Offset-Element LOE 432
Blur-ID LBI 396
Blur-Element LBE 432
MultiDoc Batch-label MB 150
Find-dup-doc MF 150
OneDoc Repeat OR 150
QA OQ 180
Extract OE 150
Table 2: Statistics of LIFBench.
@@ 3.1 Problem Definition
In this setup, the scenario description $D$ provides task background at the beginning of the prompt, and all tasks within a scenario share the same $D$. The context $X$, as the main body of the prompt, provides essential information and varies by scenario. For example, in the List scenario, $X$ is a long list, while in the OneDoc scenario, $X$ represents a lengthy processed document. The parameter $len$ represents the number of tokens in $X$. The instruction $I$ is placed at the end of the prompt, which consists of two components: (1) the instruction template ($tpl$), which outlines the task requirements, and (2) the instruction variable ($var$), representing variable part within the template. Generally, in LIFBench, $D$ and $I$ tend to be short, while $X$ is a long text with thousands of tokens.
@@ 3.2 Dataset Construction
To accurately simulate real-world application scenarios for LLMs in long-text processing, we construct 3 scenarios and 11 corresponding tasks (see Table 2), enabling a comprehensive and precise evaluation of instruction-following capabilities.
List   The List scenario is designed to assess the ability of LLMs to handle structured batches of short-item information, as seen in applications like retrieval, recommendation, and structured data processing. The context $X$ is a lengthy ordered list, and LLMs must identify elements that meet the specified criteria provided in the instructions.
@@
MultiDoc   The MultiDoc scenario tests LLMs on tasks involving coarser-grained information, such as multi-document summarization, clustering, and retrieval-augmented generation (RAG). Unlike the List scenario, the MultiDoc tasks require models to compare multiple documents, identify differences, and execute batch operations, demanding holistic attention to the input rather than focusing on a few key elements.
To construct context $X$, we curate and process datasets from diverse domains, combining them into multi-document collections with high diversity. There are two tasks in this scenario. Find-dup-doc requires LLMs to identify and output multiple duplicate documents in a specified format. In batch-label, models label each document based on specified attributes.
OneDoc   The OneDoc scenario simulates tasks involving processing of a single and very long document, typically occurring in use cases like information extraction and question answering. In this scenario, LLMs are required to identify and process the most relevant information from a vast amount of text, testing their ability to resist noise and focus on task-specific instructions.
To create this scenario, we construct an extra-long document by concatenating entries from the Paul Graham Essays dataset [^2] https://huggingface.co/datasets/sgoel9/paul_graham_essays, and tagging some sentences as key information. Then three tasks are designed: The Repeat task requires models to repeat a specified amount of key information. The Extract task instructs models to extract specific types of key information. The QA task presents models with a sentence and asks it to identify whether the sentence is labeled as key information.
@@ 3.3 Data Extension
In all three scenarios, we adjust the length of the prompt by controlling the context token count $l$. Specifically, we modify the number of elements in the List or the number of documents in MultiDoc and OneDoc to achieve the desired length. We have pre-constructed ample corpus for each scenario, supporting expansions up to 2M tokens in one prompt.
Expression   In real-world contexts, due to personality and individual differences, individuals often provide significantly varied descriptions of the same subject, which undoubtedly introduces challenges to the stability of large models in following instructions. To assess LLM robustness in this regard, we diversify instruction templates to create multiple expressions style with differing wording and syntax.
Our approach follows a four-step process, namely "Rewriting-Encoding-Clustering-Sampling (RECS)". First, we use GPT-4 Achiam et al. (2023) and Claude Anthropic (2024) to generate 40+ rewrites of each original instruction template. Next, the rewritten templates are encoded into vector representations for clustering, with the number of clusters set to the target number of rewritten instructions for each task. For instance, if a task needs five rewritten instructions, we will create five clusters. Finally, we sample from each cluster, selecting the usable template closest to the cluster center as the final diversified expression. More implementation details and evidence of effectiveness can be found in Appendix B.
Variable   Some placeholders are preset in the instruction templates to indicate variable parts of instructions, i.e., instruction variables. These variables encompass elements such as query keys (for retrieval tasks), categorization criteria (for classification tasks), and format requirements. By analyzing LLM performance across varying instruction variables, we can evaluate the models’ understanding and consistency in executing instructions. Ideally, models with strong instruction-following abilities execute instructions stably across variable conditions, while weaker models may exhibit inconsistency. Each instruction variable has been manually assigned by specific sampling spaces and rules, enabling the generation of a comprehensive instruction set. For specific examples, please refer to Appendix C.
@@
Figure 2: Rubric score distribution across different scenarios and capabilities. We carefully designed the rubric to ensure scores are as evenly distributed as possible across different scenarios and capabilities.
@@ 4.1 Automated Rubric-based Scoring
To evaluate the output quality of LLMs on LIFBench tasks, we introduce Automated Rubric-based Scoring (ARS), an accurate and efficient programmatic evaluation method. As shown in Figure 1 and Figure 2, we manually craft scoring rubrics $\mathcal{R}$ for each task. Each rubric consists of several scoring points $s$ and is assigned a weight $\tilde{s}$ according to its importance and complexity. In other words, a larger $\tilde{s}$ means a more complex task, requiring more evaluative consideration. All of these points can be assessed automatically through a program. The scoring process on task $t$ can be shown as follows:
@@
In this equation, $\mathcal{A}_{t}$ represents the model’s responses on task $t$, and $\tilde{R}_{t}$ is the sum of weights across all scoring points. The function $f_{s}(\cdot)\in[0,\tilde{s}]$ represents the average score of all outputs for scoring point $s$, as determined by the automated evaluation program. Naturally, defining $\mathcal{T}=\{t_{i}\}_{i=1}^{N_{t}}$ as the set of all tasks, the overall test result for any model on LIFBench is the weighted average of scores across $\mathcal{T}$:
@@ 4.2 Score-Capability Mapping
where $\mathbb{I}(s,c)$ is the indicator function that identifies whether a scoring point $s$ is related to the capability $c$. We manually crafte this mapping so that when a scoring point $s$ is relevant to the capability $c$, the indicator $\mathbb{I}(s,c)$ equals 1, and 0 otherwise.
@@ 4.3 Instruction Following Stability
Finally, the Instruction Following Stability (IFS) on task $t$ is calculated as the variance (denoted by $\sigma(\cdot)$) of the performance values $Y$ divided by their mean, formally expressed as:
@@ 5.1 Experiment Setup
Experiments are conducted across six context length intervals from 4k to 128k tokens, with task-specific output limits to ensure sufficient space for model generation. Token counts are calculated using GPT-4’s tokenizer [^3] https://github.com/openai/tiktoken, and truncation is applied to adjust context $X$ for models unable to process the longest contexts. Further details about the experimental setup can be found in the Appendix F.
@@
Table 4: Instruction Following Stability from different perspectives. The integers in parentheses represent rankings, where smaller numbers indicate better performance. The overall order in each partition depends on the average IFS rankings from three perspectives. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
@@ 5.2 Results on LIFBench
Task-categorized Performance   As shown in Table 3, two closed-source models achieve the highest scores, with the top score reaching only 0.758. This highlights substantial room for improvement in instruction-following capabilities. Generally, larger parameter sizes are associated with better performance. However, models with supervised instruction tuning often outperform their base counterparts, even when smaller in size. For instance, InternLM2.5-7b-chat-1m scores 0.101 higher than Llama-3.1-70B, demonstrating that fine-tuning on instruction or conversational data significantly enhances performance. While closed-source models dominate most tasks, open-source models excel in only a few (e.g., Blur-ID and Batch-label). Overall, open-source small and medium models still lag behind closed-source large models.
Capability-categorized Performance   As shown in Figure 3, closed-source models lead in all dimensions. Within each model series, the solid lines consistently enclose their corresponding dashed lines, highlighting the significant impact of instruction fine-tuning on overall performance. Notably, the Format dimension exhibits the most compact clustering, suggesting that formatting requirements are well-covered across training corpora for most models. In contrast, the Recog dimension shows marked differences, likely reflecting varying levels of effort in incorporating numerical cognition data during training.

### v3

@@ Abstract
Xiaodong Wu  Minhao Wang  Yichen Liu  Xiaoming Shi  He Yan §  Xiangju Lu §  Junmin Zhu §  Wei Zhang †‡thanks: Corresponding author. East China Normal University  §iQIYI Inc  Shanghai Innovation Institute {51255901079,51275901104,51275901148}@stu.ecnu.edu.cn  xmshi@ir.hit.edu.cn §{yanhe, luxiangju, zhujunmin}@qiyi.com  *zhangwei.thu2011@gmail.com
@@ 1 Introduction
As Large Language Models (LLMs) continue to make significant strides across practical applications Achiam et al. (2023); Chowdhery et al. (2023); Brown (2020), their performance in natural language processing (NLP) tasks has reached unprecedented levels. These tasks span text generation Que et al. (2024); Tan et al. (2024); Zhang et al. (2024b), complex reasoning Parmar et al. (2024); Chen et al. (2025), and problem-solving Lu et al. (2024); Li et al. (2024b). Despite these achievements, significant challenges remain. On the one hand, LLMs often struggle to accurately and consistently follow human instructions, such as restating input content precisely or adhering stably to specific formatting constraints He et al. (2024). On the other hand, studies show that as input length increases, the LLMs’ performance in tasks such as reasoning Levy et al. (2024), retrieval Li et al. (2024a), and general NLP Bai et al. (2023) deteriorates. These challenges pose substantial barriers to their effectiveness in real-world applications.
@@
For evaluation, traditional metrics for downstream tasks are often unsuitable for complex instruction-following scenarios Honovich et al. (2023). Moreover, while many studies rely on GPT-4 for automated and open-ended assessment, these approaches encounter limitations due to notable gaps between GPT-4 and human experts Qin et al. (2024); Jiang et al. (2023), as well as potential bias problems Wang et al. (2023). To address these challenges, we propose LIFEval, a systematic and efficient evaluation method for complex LLM responses, without relying on LLMs or human evaluators. Specifically, by designing task-specific scoring rubrics, we decompose evaluations into fine-grained and quantifiable scoring points, each of which can be assessed automatically. In addition, through the score-capability map and a novel metric—IFS, LIFEval provides insights into models’ fundamental capabilities and stability from various perspectives.
@@ 2 Related Work
Several studies focus on LLMs’ performance in long contexts. These benchmarks collect data from traditional NLP tasks (e.g., summarization and question answering (QA)) to form comprehensive datasets containing long data Shaham et al. (2022, 2023); Dong et al. (2024); Li et al. (2024c); Gavin et al. (2024). Additionally, due to the excellent performance of LLMs on open-ended tasks, some benchmarks have also designed synthetic tasks to better probe the models’ ability in math, reasoning, and logic Kwan et al. (2023); Zhang et al. (2024c); Wang et al. (2024); Li et al. (2024a); Chen et al. (2024b). For evaluation, some works adopt regular metrics (e.g., Acc, ROUGE, and BLEU), which can be obtained by automated calculations. However, some open-ended tasks cannot be effectively evaluated using these metrics, hence powerful LLMs, such as GPT-4, are used as alternative evaluators. For example, the studies An et al. (2023); Li et al. (2023) feed the model predictions and ground-truth answers to the GPT-4 evaluator, which is tasked with conducting a direct scoring or comparison to evaluate baselines’ performance on partial tasks. Unlike LIFBench, these benchmarks only assess problem-solving capabilities in long-context scenarios, overlooking challenges of complex instruction following that arise in real-world applications.
Given the complexities of evaluating instruction-following abilities, several studies introduce meaningful innovations in their evaluation methodologies to better tackle this multifaceted challenge. The studies Cook et al. (2024); Wen et al. (2024); Qin et al. (2024); Zhang et al. (2024a) decompose instructions into checklists composed of YES/NO questions or PASS/FAIL criteria, which answers should meet. CELLO He et al. (2024) defines four criteria that should be assessed as they can encompass common errors made by LLMs, and develops automated evaluation metrics to reflect models’ ability in complex instruction-following scenarios. However, these studies fail to concern evaluation in long-context instruction-following scenarios and largely overlook stability in the instruction-following process. Additionally, while the study Sakai et al. (2024) explores the impact of various prompt templates and languages on LLM, it focuses solely on NLU tasks.
In summary, existing benchmarks either emphasize long-context scenarios or instruction-following capabilities, whereas LIFBench uniquely targets both simultaneously.
@@ 3.1 Problem Definition
In this setup, the scenario description $D$ provides task background at the beginning of the prompt, and all tasks within a scenario share the same $D$. The context $X$, as the main body of the prompt, provides essential information and varies by scenario. For example, in the List scenario (three scenarios are constructed in LIFBench, see Section 3.2), $X$ is a long list, while in the OneDoc scenario, $X$ represents a lengthy processed document. The parameter $len$ represents the number of tokens in $X$. The instruction $I$ is placed at the end of the prompt, which consists of two components: (1) the instruction template ($tpl$), outlining the task requirements, and (2) the instruction variable ($var$), representing variable part within the template. Generally, in LIFBench, $D$ and $I$ tend to be short, while $X$ is a long text with thousands of tokens.
@@ 3.2 Dataset Construction
To simulate real-world LLM applications in long-text processing, we construct 3 scenarios and 11 tasks (see Table 2) based on the following principles: (1) Task Diversity: Tasks should encompass varied constraints (e.g., format, quantity) to evaluate different instruction-following abilities; (2) Performance Distinguishability: Tasks must balance simplicity and complexity to distinguish model performance; (3) Input Scalability: Tasks should support extended input lengths to assess long-context capabilities; (4) Automated Evaluation: Task constraints can be assessed through an automated program.
List   The List scenario tests how well LLMs can handle structured lists, such as retrieving specific items and processing structured data. The input $X$ is an ordered list with UUIDs and natural language instructions.
@@
MultiDoc   The MultiDoc scenario evaluates how well LLMs process multiple documents, such as summarization and retrieval. Models need to compare documents, find differences, and handle batch operations.
The input $X$ consists of multi-document collections from diverse sources. Each document has six fields: "text", "id", "iD2", "title", "date", and "source", with a length of 300–500 tokens. Tasks include Find-dup-doc, which finds duplicate documents, and Batch-label, which assigns labels to documents based on given attributes.
OneDoc   The OneDoc scenario tests how well LLMs process a single long document, such as extracting key information or answering questions.
A long document is created by combining essays from the Paul Graham Essays dataset [^2] https://huggingface.co/datasets/sgoel9/paul_graham_essays. Some sentences are marked as key information. Tasks include Repeat, where models repeat a given amount of key information; Extract, where models extract specific key details; and QA, where models check if a sentence is labeled as key information.
@@ 3.3 Data Extension
In all three scenarios, we adjust the length of the prompt by controlling the context token count $l$. Specifically, we modify the number of elements in the List or the number of documents in MultiDoc and OneDoc to achieve the desired length. Ample corpus are pre-constructed for each scenario, supporting expansions up to 2M tokens in one prompt.
Expression   In real-world contexts, due to personality and individual differences, individuals often provide significantly varied descriptions of the same subject, which undoubtedly challenges the stability of large models in following instructions. To assess LLM robustness in this regard, we diversify instruction templates to create multiple expressions style with differing wording and syntax.
Our approach follows a four-step process, namely "Rewriting-Encoding-Clustering-Sampling (RECS)". First, to ensure diversity and mitigate biases from any single model, we use GPT-4 Achiam et al. (2023) and Claude Anthropic (2024) to generate 40+ rewrites of each original instruction template, complemented by a subsequent manual review to further validate and refine the outputs. Next, the rewritten templates are encoded into vector representations for clustering, with the number of clusters set to the target number of rewritten instructions for each task. For instance, if a task needs five rewritten instructions, we will create five clusters. Finally, from each cluster, we select the usable template nearest to the center as the final diversified expression. Further details and effectiveness evidence are provided in Appendix B.
  Scenario Task ID #Exp. #Var. #Data
List Single-ID LSI 5 6 180
Multi-ID LMI 5 5 150
Offset-ID LOI 11 6 396
Offset-Element LOE 12 6 432
Blur-ID LBI 11 6 396
Blur-Element LBE 12 6 432
MultiDoc Batch-label MB 5 5 150
Find-dup-doc MF 5 5 150
OneDoc Repeat OR 5 5 150
QA OQ 5 6 180
Extract OE 5 5 150
Table 2: Statistics of LIFBench. #Exp. and #Var. represent the count in the data extension of Expression and Variable.
Variable   Some placeholders are preset in the instruction templates to indicate variable parts of instructions, i.e., instruction variables. These variables encompass elements such as query keys (for retrieval tasks), categorization criteria (for classification tasks), and format requirements. For position-related or numerical variables, we maintain an even distribution, and manually adjust other variables during task iterations to precisely control the task difficulty (see examples in Appendix C). By analyzing LLM performance across varying instruction variables, we can evaluate the models’ understanding and consistency in executing instructions. Ideally, models with strong instruction-following abilities execute instructions stably across variable conditions, while weaker models may exhibit inconsistency.
@@
Figure 2: Rubric score distribution across different scenarios and capabilities. Scores across different scenarios and capabilities tend to be evenly distributed.
@@ 4.1 Automated Rubric-based Scoring
To evaluate the output quality of LLMs on LIFBench tasks, we introduce Automated Rubric-based Scoring (ARS), an accurate and efficient programmatic evaluation method. As shown in Table 6 in Appendix D, we manually craft scoring rubrics $\mathcal{R}$ for each task. Each rubric consists of several scoring points $s$ and is assigned a weight $\tilde{s}$ according to its complexity and difficulty. In other words, a larger $\tilde{s}$ means greater complexity, requiring more steps for evaluation. All of these points can be assessed automatically through a program. The scoring process on task $t$ can be shown as follows:
@@
In this equation, $\mathcal{A}_{t}$ represents the model’s responses on task $t$, and $\tilde{R}_{t}$ is the sum of weights across all scoring points. The function $f_{s}(\cdot)\in[0,\tilde{s}]$ represents the average score of all outputs for scoring point $s$. Its implementation relies on programmatic evaluation pipelines tailored to each scoring point. For example, in structural verification (e.g., JSON Dict format, see Program 1), we use a multi-stage detection process that includes symbol check, Parsing check, and KV Check. These stages progressively impose stricter requirements on the model’s output, with final scores based on performance at each stage. Overall, while the core logic is manually engineered according to the rubric, the execution is fully automated through programs. We provide detailed considerations regarding the scoring rubric, as well as additional examples of evaluation programs, in Appendix D.
Naturally, defining $\mathcal{T}=\{t_{i}\}_{i=1}^{N_{t}}$ as the set of all tasks, the overall test result for any model on LIFBench is the weighted average of scores across $\mathcal{T}$:
@@
Although a well-designed scoring rubric enables LIFEval to provide a more comprehensive and rigorous evaluation, the relatively high human cost inevitably limits its generalizability. To address this, when extending LIFBench with new samples or adapting LIFEval to other datasets, users may adopt a simplified approach by replacing weighted averaging in Equation 2 and 3 with a simple mean. This streamlined method bypasses the rubric design process while still satisfying the requirements of many practical applications.
@@ 4.2 Score-Capability Mapping
where $\mathbb{I}(s,c)$ is the indicator function that identifies whether a scoring point $s$ is related to the capability $c$. We manually crafted this mapping so that when a scoring point $s$ is relevant to the capability $c$, the indicator $\mathbb{I}(s,c)$ equals 1, and 0 otherwise. Table 6 provides more detailed information on this. Importantly, as shown in Figure 2, scores across different scenarios and capabilities are carefully balanced during data collection to ensure less bias in LIFEval.
@@ 4.3 Instruction Following Stability
Finally, the Instruction Following Stability (IFS) on task $t$ is calculated as the standard deviation (denoted by $\sigma(\cdot)$) of the performance values $Y$ divided by their mean, formally expressed as:
@@ 5.1 Experiment Setup
The experiments were conducted across six context length intervals, ranging from 4k to 128k tokens, with task-specific output limits to ensure sufficient space for model generation. Token counts are calculated using GPT-4’s tokenizer [^3] https://github.com/openai/tiktoken, and truncation is applied to adjust context $X$ for models unable to process the longest contexts. In data extension, 5–6 template variants are created for each original instruction, and each instruction variable has 5–10 candidates for sampling. Further details can be found in Appendix E.
@@
Table 4: Instruction Following Stability (IFS) from three perspectives, with rankings in parentheses (smaller is better). The overall order is based on average IFS rankings. $\dagger$ denotes truncation of context $X$ on the longest interval.
@@ 5.2 Results on LIFBench
Task-categorized Performance   As shown in Table 3, two closed-source models achieve the highest scores, with the top score reaching only 0.758. This highlights substantial room for improvement in instruction-following capabilities. Generally, larger parameter sizes are associated with better performance. However, models with supervised instruction tuning often outperform their base counterparts, even when smaller in size. For instance, InternLM2.5-7b-chat-1m scores 0.101 higher than Llama-3.1-70B, demonstrating that fine-tuning on instruction or conversational data significantly enhances performance. While closed-source models dominate most tasks, open-source models excel in only a few (e.g., Blur-ID and Batch-label).
Moreover, compared to LSI, most of models perform worse on LOI and LBI tasks, likely due to LSI’s closer alignment with their training data, highlighting the need to enhance their ability to handle complex instructions. Similarly, the model performance on LBE (LOE) tasks is better than that on LBI (LOI), indicating that referencing specific elements is easier than using non-semantic IDs.
Capability-categorized Performance   As shown in Figure 3, closed-source models lead in all dimensions. Within each model series, the solid lines consistently enclose their corresponding dashed lines, highlighting the significant impact of instruction fine-tuning on overall performance. Notably, the Format dimension shows the tightest clustering, suggesting that formatting requirements are well-covered across training corpora for most models. In contrast, the Recog dimension shows marked differences, likely reflecting varying levels of effort in incorporating numerical cognition data during training.
@@
Figure 6: The impact of increasing context length on ARS scores in tasks with different instruction complexities.